OpenStudy (anonymous):

As waves enter shallow water and decrease in speed, wave height increases and eventually a wave will break. The point at which a wave will break can be calculated using the formula for wave steepness: steepness = wave height/wavelength. When the steepness of a wave reaches 1/7, the wave will break. If the wavelength of a wave is 50 m, at what height will the wave break?

OpenStudy (anonymous):

This thing set itself up. \[Steepness = \frac{ height }{ wavelength }\] They tell us that the steepess is 1/7 and the wavelength is 50m. It's just a matter of solving this: \[\frac{ 1 }{ 7 } = \frac{ height }{ 50 }\] Cross multiply and solve.